At a picnic a cooler filled with ice there are 100 cans of soda. With a mixture of sodas. 30 cans of cola, 40 cans of orange, 20 cans of root beer and 10 cans of cream soda. What are the odds of, if you reach in and you select just ONE can, and that can is either a cola or a orange soda?
This is what I have done so far.
N= 100, r= 30
= 30/100 = 3/10 =0.3
n=100, r=40
= 40/100= 4/10 = 0.4
= 0.3 + 0.4 = 0.7 times chance of getting either a cola or a can of orange soda.
-------------------------------------------------------------------------------
Please help me!

Answer by edjones(8007)

P=70/100=7/10 probability of success.
3/10 probability of failure.
Odds in favor of success are 7/10 to 3/10 or simply 7 to 3.
.
Ed